OpenStudy (juana02):

Does the probability of rolling a sum of 2 appear to be the same as the probability of rolling a sum of 7? Explain your answer.

OpenStudy (mimi_x3):

Empirical or theoretical probability?

OpenStudy (mimi_x3):

I have seen your question yesterday and If I remember correctly the question was referring to empirical probability It was basically based on your observations

OpenStudy (mimi_x3):

You had to roll the die 60 times and then record the amount of times it landed on each number

OpenStudy (juana02):

yes but this question is different im having trouble with probablity

OpenStudy (mimi_x3):

ok well then lets pretend or assume that this is theoretical probability Ok when you roll a die once there are a total of 6 diff options - it can land on any of 6 number. It can land on 1 , 2, 3, 4, 5, and 6 The probability that it will land on 1 is ... \[ \text{ Probability of 2} = \frac{\text{ How many times it lands on 2}}{\text{ Total # of outcomes}}\] Now lets look above ... How many outcomes are there??? 6 . It can land on 6 numbers ... 1,2,3,4,5 and 6 How many times does it land on a 2??? Only once so 1 Therefore the probability that it will land on 2 is \[ \text{ Probability of 2} = \frac{\text{ 1}}{\text{ 6}}\]

OpenStudy (mimi_x3):

Oh whoopppsssss Didnt read the question ... it asks for sum of 2 and sum of 7

OpenStudy (mimi_x3):

ok this all depends on how many times you roll a die

OpenStudy (mimi_x3):

Is there more to the question?

OpenStudy (juana02):

well i have more questions

OpenStudy (mimi_x3):

ya but is there a first part to the question?

OpenStudy (mimi_x3):

Like im assuming that you havta roll the die twice

OpenStudy (juana02):

so for example there is not a single answer? your answer will be differfent then mine depending on how many times i roll the dice?

OpenStudy (juana02):

no thats the whole question

OpenStudy (mimi_x3):

Not really ... thats only in empirical probability

OpenStudy (mimi_x3):

In theoretical probability its always the same its always a 1/6 that you will land on a 2

OpenStudy (mimi_x3):

ok so lets see when you roll the die twice The only way you can get a sum of 2 is if you roll a 1 on the first time and a 1 on the second time you roll the die 1+1=2

OpenStudy (mimi_x3):

So the probability that you will get a 1 is a 1/6 so (1/6)*(1/6)=1/36 The probability that you will roll on 1 twice in a row is 1/36

OpenStudy (mimi_x3):

for a sum of 7 you can need to roll on a 3 and 4 so its 3+4=7 Now there are 2 ways to get a sum of 7 and I will show you below Option#1 Roll #1 we get 3 Roll #2 we get 4 Probability that will land on 3 is a 1/6 Probability that will land on 4 is a 1/6 So probability that will first land on 3 and then 4 is (1/6)*(1/6)=1/36 Option#2 Roll #1 we get 4 Roll #2 we get 3 Probability that we will land on 4 is 1/36 Probability that will land on 3 is 1/6 Probability that will land on both is (1/6)(1/6)=1/36 The probability that we will get either option 1 or option 2 is (1/36)+(1/36)=2/36 So therefore the probability that will get a sum of 7 is 2/36

OpenStudy (mimi_x3):

So does the probability appear to be the same?????

OpenStudy (mimi_x3):

Ok i dont think i explained this clearly ...

OpenStudy (juana02):

yes

OpenStudy (mimi_x3):

hmmm i gotta go study but maybe @freckles would like to help :P

OpenStudy (freckles):

So how many dice do you have? 2,3,4,5? I assumed it was more than 1 since you said sum of the die.

OpenStudy (freckles):

If it is 2 die:  This picture represents all the possible outcomes of rolling two fair die together